the new seating chart in Mr. Lian's class of 26 students was randomly generated. What is the probability that Jamila, Candace, and Haley are in the first, second, and third seats respectively?
Answer by stanbon(75887)

the new seating chart in Mr. Lian's class of 26 students was randomly generated. What is the probability that Jamila, Candace, and Haley are in the first, second, and third seats respectively?
--------------
# of ways to succeed: 1
# of possible arrangements: 26P3 = 26*25*24 = 15600
--------------------
Probability of that arrangement = 1/15600
======================
Cheers,
Stan H.